alones (alones):

Solve the following quadratic equation using the quadratic formula. Which of the following expressions gives the numerators of the solutions?

alones (alones):

\[10x^2 - 19x + 6 = 0\]

OpenStudy (kittiwitti1):

a. \(\text{What "following expressions"?}\) b. The quadratic formula for variable \(x\) is the following: \(x=\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}\). Note that the discriminant is under the radical... \(\sqrt{b^{2}-4ac}\)

OpenStudy (kittiwitti1):

so for this equation: a = 10 b = -19 c = 6\[x=\frac{-(-19)\pm\sqrt{(-19)^{2}-4(10)(6)}}{2(10)}\]

OpenStudy (kittiwitti1):

simplified\(\rightarrow\)\[\frac{19\pm\sqrt{361-240}}{20}\]

alones (alones):

Oka thanks, let me solve that :3

alones (alones):

Okay i got 19 ± 11

OpenStudy (johnweldon1993):

*Over 20! Don't forget that :)* \[\large \frac{19 \pm 11}{20}\]

OpenStudy (kittiwitti1):

But they just want the numerator, right? @johnweldon1993

OpenStudy (johnweldon1993):

Reasons why I should read the full question first >.< Thank you @kittiwitti1 :)

OpenStudy (kittiwitti1):

lol P: no problem

OpenStudy (anonymous):

Refer to the attachment from Mathematica.